1: the following are taking from national income accounts of a country

A: GDP=600 B: gross investment   800 C; net investment 200

D: consumption 4000 E: govt . purchases   1100 F: govt. 30(tax revenue-govt . spending)

find out: I: NDP II: disposable income III: net export IV: personal saving
Answer #1

NDP = GDP - Depreciation

Depreciation = Cross Investment - Net Investment

. = 800 - 200

= 600

NDP = GDP - Depreciation

= 6000 - 600

= 5400

DISPOSAL INCOME = GDP - Purchases - Govt Surplus

. = 6000 - 1100 + 30

= 6000 - 1130

= 4870

NET EXPORTS = GDP - CONSUMPTION - INVESTMENT - Govt Purchases

. = 6000 - 4000 - 800 - 1100

= 100

PERSONAL SAVING = DISPOSAL INCOME - CONSUMPTION

. = 4870 - 4000

= 870